OpenStudy (anonymous):

2An airplane initially at rest begins to accelerate along the landing strip and will detach from the strip after t1 = 15s acceleration at a distance s1 = 600m from its initial position. Find the acceleration a (assume to be uniform) of the plane.

OpenStudy (anonymous):

Again using newtons second law equation s=ut+0.5a(t^2) 600=0.5*a*(15)^2 a=1200/225 m/s^2 a=5.33m/s^2. u=0 as the plane is initially at rest.

OpenStudy (anonymous):

am i right @ hashmaz

OpenStudy (anonymous):

Yes u are right.